A cleaning company charges a fixed amount for a house call and a second amount for each room it cleans. The total cost to clean 6 rooms is $250 and the total cost to clean 8 rooms is $320. How much would this company charge to clean two rooms?
Answer by lwsshak3(11628)

A cleaning company charges a fixed amount for a house call and a second amount for each room it cleans. The total cost to clean 6 rooms is $250 and the total cost to clean 8 rooms is $320. How much would this company charge to clean two rooms?
***
let y=fixed amt for each house call
let x=cost for each house it cleans
..
y+6x=250
y+8x=320
subtract:
-2x=-70
x=35
y=250-6x=250-210=40
How much would this company charge to clean two rooms=40+2*35=40+70=$110